A $1.60 assortment of cards contains $0.20 cards and $0.15 cards. The number of $0.20 cards is 2 less than half the number of $0.15 cards. Find the number of each type of card.
Thank you!!

Answer by Paul(988)

Let the number of $0.15 be x
Let the number of $0.20 be 0.5x-2
Expression:
0.20(0.5x-2)+0.15x=1.60
0.1x-0.4+0.15x=1.60
0.25x=2
x=8
0.5(8)-2 = 2

Hence, there were 8 $0.15 cards and 2 $0.20 cards.
Paul.